OpenStudy (chrismatias):

The solution set for the inequality x + 4 (x - 3) includes -11 as an element. true or False?

jimthompson5910 (jim_thompson5910):

It depends on what is between the `x+4` and the `(x-3)`. The inequality sign isn't showing up.

OpenStudy (chrismatias):

x + 4>*or equal to* 1/2

jimthompson5910 (jim_thompson5910):

1/2 is on the right side? not x-3 ?

OpenStudy (chrismatias):

1/2 on the left with (x+4)

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the entire problem?

OpenStudy (chrismatias):

The solution set for the inequality x + 4>*or equal to* 1/2 (x - 3) includes -11 as an element.

jimthompson5910 (jim_thompson5910):

So the inequality is this right? \[\Large x+4 \ge \frac{1}{2}(x-3)\]

jimthompson5910 (jim_thompson5910):

I'm going to assume it is

jimthompson5910 (jim_thompson5910):

Replace EVERY copy of 'x' with -11. Then simplify \[\Large x+4 \ge \frac{1}{2}(x-3)\] \[\Large -11+4 \ge \frac{1}{2}(-11-3)\] \[\Large -7 \ge \frac{1}{2}(-14)\] \[\Large -7 \ge -7\] Since \[\Large -7 \ge -7\] is a true inequality, this means the original inequality is true when x = -11. So -11 is definitely an element of the solution set.